OpenStudy (anonymous):

Simplify the expression: -4(yz+2)-6yz+1+y^2

OpenStudy (uri):

@MarthaHearst Try doing this one yourself. :)

OpenStudy (uri):

I'll guide you if you get confused.

OpenStudy (anonymous):

-4yz-8-6yz+1+y^2=-10yz-7+y^2 Just apply the distributive property..

OpenStudy (anonymous):

Distributive property: A(B+C)=AB+AC

OpenStudy (anonymous):

So -4yz 2+3 ?

OpenStudy (sawpnilsrabon):

-4(yz+2)-6yz+1+y^2 -4yz-8-6yz+1+y^2 -10yz-7+y^2 y^2-10yz-7